40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 4930     Entry time: Fri Jul 1 18:41:53 2011
Author: Jamie 
Type: Update 
Category: SUS 
Subject: Core optic sus damping controllers normalized 

I found many of the core optic (ETMs, ITMs, BS, PRM, SRM) suspension DOF damping controllers (SUSPOS, SUSPIT, SUSYAW, SUSSIDE) to be in various states of disarray:

  • Many of the controllers did not have their "Cheby" and "BounceRoll" filters switched on.
  • Some of the controllers didn't even have the Cheby or BounceRoll filters at all, or had other different filters in their place.
  • ETMY was particularly screwy (I'll make a separate follow-up post about this)
  • A bunch of random other unused filters lying around.
  • oplev servos not on
  • etc.

I went around and tried to clean things up, by "normalizing" all of the DOF damping filter banks, ie. giving them all the same filters and clearing out unused filters, and then turning on all the appropriate filters in all core optic damping filter banks ("3:0.0", "Cheby", "BounceRoll").  I also went sure that all the outputs were properly on, and the oplev servos were on.

A couple of the optics had to have their gains adjusted to compensate for filter changes, but nothing too drastic.

Everything now looks good, and all optics are nicely damped.

I didn't touch the MC sus damping controllers, but they're in a similar state of disarray and could use a once-over as well.

 

ELOG V3.1.3-